Cher opened a bakery that manufactures packaged Cupcakes in LA. Using the Plantwide allocation system, they have determined that the cost per cupcake is $1.30. Since operating expenses in California are high, Cher wants a gross profit percentage of 75%. For what price should they sell their product? Round answer to the nearest penny (e.g. x.xx).
Answer #1

Solution Let the sales price be a Operating expesnes + Gross profit = Sales $1.3+(a x 75%) = a $1.3 = a -0.75a $1.3 = 0.25 a